# Data-Free Class Incremental Gesture Recognition via Synthetic Feature Sampling

### v1

@@ Abstract
Data-Free Class Incremental Learning (DFCIL) aims to enable models to continuously learn new classes while retaining knowledge of old classes, even when the training data for old classes is unavailable. Although explored primarily with image datasets by researchers, this study focuses on investigating DFCIL for skeleton-based gesture classification due to its significant real-world implications, particularly considering the growing prevalence of VR/AR headsets where gestures serve as the primary means of control and interaction. In this work, we made an intriguing observation: skeleton models trained with base classes(even very limited) demonstrate strong generalization capabilities to unseen classes without requiring additional training. Building on this insight, we developed Synthetic Feature Replay (SFR) that can sample synthetic features from class prototypes to replay for old classes and augment for new classes (under a few-shot setting). Our proposed method showcases significant advancements over the state-of-the-art, achieving up to 15% enhancements in mean accuracy across all steps and largely mitigating the accuracy imbalance between base classes and new classes.
@@ 1 Introduction
Class incremental learning (CIL) refers to the learning paradigm wherein models are continuously updated as new class data is introduced. This area has gained significant attention as researchers recognize that data in real-world scenarios typically arrives incrementally rather than all at once. Although considerable progress has been made in class incremental learning[51, 25, 34, 8, 41], most of the research has focused on experimenting with images as the primary modality. However, the recent emergence of large pre-trained, visual or visual-language models, such as CLIP[33], DINO[4], SAM[19] and their variations, has demonstrated remarkable generalizability, showcasing impressive zero-shot performance. This development, to some extent, challenges the significance of class incremental learning research focused solely on images. We wouldn’t be too surprised if leveraging large pre-trained vision models could surpass state-of-the-art performance on common image datasets that most class incremental learning studies tested on, either in zero-shot set or given a few shot prompts.
@@
Therefore, we argue that class incremental learning research should address specific real-world problems or be differentiated by modalities, as different modalities may exhibit distinct interclass properties as the CIL algorithm primarily experimented on image datasets might not be a universal solution adaptable to other modalities. Unlike the abundance of image sources readily available from the Internet, human-related data sources are more limited due to factors such as labor costs and privacy concerns associated with collecting data from individuals. These types of data sources may align better with the continuous availability of data. With the increasing trend in Virtual Reality (VR) technology and the growing market penetration of high-quality VR devices like Oculus Quest and Vision Pro, personal VR/AR devices are becoming more prevalent.
@@
Gesture-based interaction is the primary means of controlling AR/VR headsets [15, 16]. Class incremental learning for skeleton-based gesture recognition shows promising value; for example, users may register and customize their own gestures for various functions. However, data privacy and security are significant concerns for human-related computer vision tasks, as even skeletal gestures can inadvertently reveal personal identity information. Thus, restricting the setting to be data-free, meaning that only data used to train newly introduced classes is available without access to training data for old classes, becomes especially important and natural.
@@
In this study, we explore Data Free Class Incremental Learning(DFCIL) for Skeleton Gesture, a new research domain introduced by BOAT-MI[1]. Unlike BOAT-MI which involves data synthesis via model inversion and retraining of the whole model. We decide to keep the feature extractor consistently frozen, as we observe that the model trained for skeletal data exhibits a much greater ability to generalize to new classes than images without additional updating. Then, by modeling class feature space as a multivariate normal distribution characterized by a prototype defined by its mean and covariance, we can sample synthetic features that can be used for both old class replay and new class augmentation.
@@
Based on this observation, we proposed a novel Synthetic Feature Replay (SFR) algorithm that circumvents the need for challenging synthetic data generation, which is typically used in existing DFCIL methods. Our approach is not only faster and simpler, but also significantly outperforms current SOTA methods across all tested datasets.
Additionally, we examined the effect of shot size within the Few Shot Class Incremental Learning (FSCIL) context and found that sampling synthetic features for new classes combined with a carefully designed filtering mechanism can alleviate performance degradation caused by limited data availability.
@@ 2 Related Work
FSCIL [41] is a setup that imposes additional restrictions on the sample size of incremental classes beyond those in DFCIL. One strategy widely used in few-shot class incremental learning is to keep the backbone encoder frozen and update only the classifier to adapt to new classes. These methods focus either on creating better pre-trained base models [56, 40, 57], or designing novel classifiers [53]. Constructing a class prototype [27, 14, 3, 21], originally proposed by ProtoNet [39], primarily used for a few-shot learning, has also proved useful in creating the classifier for FSCIL as well [58, 45]. Our approach also adopts such a frozen encoder strategy, but instead of directly using class prototypes to build the classifier, we sample synthetic features from class prototypes and train the classifier along with the sampled features.
@@
Using skeletal data, instead of raw image frames, for gesture recognition, shows great value for its reduced data complexity and therefore effective processing and privacy protection. Deep learning approaches such as CNN-LSTM [31], GCNs[49, 7, 23], transformers [6, 36, 55, 44], typically show superior performance for skeleton-based gesture recognition. ST-GCN[49] is a popular GCN-based model that can automatically learn both spatial and temporal patterns. DG-STA [6] is a transformer-based model that applies multi-head attention to spatial and temporal edges, respectively, for hand gesture recognition.
Class incremental learning for skeleton data, especially in the data-free setting, is still a valuable but underexplored domain. To the best of our knowledge, BOAT-MI [1] represents pioneering work in investigating data-free class incremental learning of gestures by developing a model inversion-based approach to generate synthetic samples and replay them during training new classes. Therefore, we follow their setup to run our experiments and benchmark against them.
@@ 3.2 Metric Learning
3.3 Instantaneous Forgetting Measure (IFM)
The central dilemma in class incremental learning revolves around balancing the preservation of knowledge from old tasks with the ongoing learning of new tasks. However, many existing studies[26, 53, 57, 8] tend to overprioritize overall accuracy as the primary evaluation metric, often overlooking the imbalance between the accuracy of the base classes and the incremental classes. For instance, training directly on new classes without any form of regularization can lead the model to quickly forget old knowledge, causing the new class accuracy to significantly outperform the old classes. On the contrary, frozen encoder-based image CIL approaches tend to perform much poorer on newly introduced classes compared to base classes that are inherently trained for the base model.
@@ 3.3 Instantaneous Forgetting Measure (IFM)
where represents the local accuracy solely over the novel classes introduced during a continual learning step, while signifies the global accuracy over all classes encountered by the model up to that stage, including the novel classes. IFM is expressed as a percentage, with 0% indicating perfect balance () and 100% indicating pure imbalance (either or ).
4 What to Replay? Embeddings or Samples
Synthetic replay-based DFCIL approaches[8, 37, 52, 13, 1] typically involve training an additional generator or using model inversion to create fake data. However, generating such samples poses significant challenges, as pre-training a generator during the previous step can be cumbersome, and the quality of samples generated through model inversion may suffer, especially for skeletal gesture movements that entail both spatial and temporal changes. BOAT-MI[1] proposed a novel boundary-aware sampling mechanism from class prototypes to generate the best synthetic samples via model inversion, but they still have trouble generating high-quality samples.
Thus, is there a way to bypass the challenging data synthesis step? Our proposed approach directly utilizes synthetic features sampled from prototypes to replay the old class and facilitate the training of the new class, when the samples of the newly class are limited, as illustrated in Figure 1. The effectiveness of our approach is rooted in a key observation.
@@ 4 What to Replay? Embeddings or Samples
However, we found that using a frozen encoder is particularly well suited for skeleton modality, as our experiments unveiled an interesting discovery: models trained on skeletal-based gesture movement data can automatically generalize well to new, unseen classes without additional training. This enables the formation of compact feature spaces that establish distinguishable boundaries from old and new classes. A simple experiment on Shrec-2017[9], as shown in Figure 3(b), illustrates how within-class features cluster together, forming compact class feature spaces with minimal overlap between different classes. This phenomenon appears to be specific to skeletal data, possibly due to its structured nature. Taking a hand gesture as an example, regardless of the specific movement (e.g., pinch or swipe), all data samples are fundamentally hand motions, and the sample structure should always adhere to the physical appearance of a hand. Thus, treating all data as a single class with various subclasses, explainable by semantic relationships, seems reasonable. Previous studies [30, 10, 22, 12, 32, 29] have shown that both large image and language models demonstrated a strong relationship between feature space and semantic information, especially within a large category. Furthermore, an observation that supports this claim is the significantly higher DFCIL performance on CUB200 (lower SAD and IFM in the last session) compared to two other image datasets, as shown in Figure 2. This can possibly be attributed to the fine-grain nature of the CUB200 which contains 200 subcategories of a single large class - bird. The skeletal encoder, even trained with very limited data, appears to be capable of capturing such a feature relationship. This observation forms the basis for our feature sampling approach.
Using the mean of class feature space as class prototypes to build classifiers is a prevalent and conventional approach in FSCIL [27, 45, 53, 58]. However, relying solely on the mean may only identify a rough location within the entire feature space and is probably not sufficient to provide an accurate approximation of class feature space. Thus, in addition to the class mean , following [1, 50], we also incorporate the covariance , providing a more comprehensive depiction of the class feature space. With the class prototype established, [50] models it as a Gaussian distribution and sample augmented features for few shot learning. Our approach also found it to be suitable for rehearsing old class data, as described in Algo. 1. By making the assumption that each class feature space follows a multivariate Gaussian distribution, we can formulate the sampling processing as:
@@
where is the synthetic features space of class sampled from the distribution, is the label space.
@@
Considering that feature representations often reside in high-dimensional spaces, a potential issue arises with the dimensionality-collapse problem. This arises when the computed class covariance fails to be positive semi-definite, contradicting the assumption of a normal distribution. One proposed solution is to employ singular value decomposition (SVD) to reduce the dimensionality of the feature space. Subsequently, sampling can be conducted in this reduced feature space, followed by reverting the sampled features back to their original dimension size, as elaborated in Algo. 1.
@@ 4.1 Effect of Supervised Contrastive learning
In contrastive learning framework, a model is trained by pulling together samples from the same classes in embedding space while pushing apart samples coming from different classes. Model trained with constraive learning are proven to generate more robust and generailzable representation, making it effective for multiple downstream tasks, and have a superior transfer ability than its cross-entropy counterpart [18]. Some previous work has shown its effectiveness in continual learning studies [11, 40].
Therefore, during the pre-training stage, we experimented with both vanilla cross-entropy loss and supervised contrastive loss. The supervised contrastive loss for a pair of samples and can be represented as:
Tables 1 and 2 show the performance when the base model is trained with Supcon and cross entropy. For Shrec-2017, which is a small dataset, the model trained with Supcon indicates a great performance for both overall accuracy and more significantly for the IFM score, which is half compared to model trained with cross entropy. However, for the much larger Ego-Gesture3D, the significance of contrastive learning is not significant.
@@ 4.3 SOTA Comparison
Method Task 0 Task 1 Task 2 Task 3 Task 4 Task 5 Task 6 Mean (Task G IFM G IFM G IFM G IFM G IFM G IFM G IFM Full-Shot 89.4 88.4 3.3 87.3 2.1 86.0 0.5 85.3 0.2 83.0 0.2 82.3 1.9 85.9 1.2 10-Shot w/o Aug 89.4 88.0 6.2 86.3 5.8 81.8 12.1 85.4 1.2 78.8 6.9 78.9 6.2 84.1 5.5 5-Shot w/o Aug 89.4 85.9 20.7 82.8 15.8 79.9 13.2 78.5 9.8 75.9 11.1 74.4 10.1 81.0 11.5 5-Shot w/ Aug 89.4 87.2 10.0 84.2 10.5 79.2 9.7 79.8 5.9 75.0 10.2 74.4 5.9 81.3 7.5
Method Task 0 Task 1 Task 2 Task 3 Task 4 Task 5 Task 6 Mean (Task G IFM G IFM G IFM G IFM G IFM G IFM G IFM Full-Shot 75.2 75.6 2.5 73.1 3.0 72.7 1.7 72.7 2.4 70.2 2.2 69.5 3.2 72.7 2.1 10-Shot w/o Aug 75.2 73.6 14.5 71.8 9.0 68.9 11.5 66.5 8.1 66.3 6.5 63.0 13.6 69.4 9.0 5-Shot w/o Aug 75.2 72.1 15.8 68.4 25.2 65.1 18.6 64.4 15.2 60.8 19.3 58.9 20.2 66.4 16.3 5-Shot w/ Aug 75.2 72.3 10.8 69.1 14.2 66.6 15.7 66.4 12.5 62.6 11.9 58.3 18.0 67.2 11.9

### v2

# Continual Gesture Learning without Data via Synthetic Feature Sampling

@@ Abstract
Data-Free Class Incremental Learning (DFCIL) aims to enable models to continuously learn new classes while retaining knowledge of old classes, even when the training data for old classes is unavailable. Although explored primarily with image datasets by researchers, this study focuses on investigating DFCIL for skeleton-based gesture classification due to its significant real-world implications, particularly considering gestures serve as one of the primary means of control and interaction in VR/AR, smart home systems and beyond. In this work, we made an intriguing observation: skeleton models trained with base classes(even very limited) demonstrate strong generalization capabilities to unseen classes without requiring additional training. Building on this insight, we developed a new model-updating paradigm - Synthetic Feature Replay (SFR) that directly synthesize fake features within the embedding space from saved class prototypes, instead of generating synthetic data, to replay for old classes and augment for new classes (under a few-shot setting). Our proposed method showcases significant advancements over the state-of-the-art, achieving up to 15% enhancements in mean accuracy across all steps and largely mitigating the accuracy imbalance between base classes and new classes.
@@ 1 Introduction
Class incremental learning (CIL) refers to the learning problem wherein models are continuously updated as new class data is introduced. This area has gained significant attention as researchers recognize that data in real-world scenarios typically arrives incrementally rather than all at once. Although considerable progress has been made in class incremental learning[50, 25, 34, 8, 41], most of the research has focused on experimenting with images as the primary modality. However, the recent emergence of large pre-trained, visual or visual-language models, such as CLIP[33], DINO[4], SAM[19] and their variations, has demonstrated remarkable generalizability, showcasing impressive zero-shot performance. This development, to some extent, challenges the significance of class incremental learning research focused solely on images.
@@
Therefore, we argue that class incremental learning research should address specific real-world problems or be differentiated by modalities, as different modalities may exhibit distinct interclass properties as the CIL algorithm primarily experimented on image datasets might not be a universal solution adaptable to other modalities. Unlike the abundance of image sources readily available from the Internet, human-related data sources are more limited due to factors such as labor costs and privacy concerns associated with collecting data from individuals. These types of data sources may align better with the continuous availability of data.
@@
Gesture-based interaction is the primary means of controlling AR/VR headsets [15, 16] and is also widely applied in various real-world scenarios, including smart homes, education, and more. Class incremental learning for skeleton-based gesture recognition shows promising value; for example, users may register and customize their own gestures for various functions. However, data privacy and security are significant concerns for human-related computer vision tasks, as even skeletal gestures can inadvertently reveal personal identity information. Thus, restricting the setting to be data-free, meaning that only data used to train newly introduced classes is available without access to training data for old classes, becomes especially important and natural.
@@
In this study, we explore Data Free Class Incremental Learning(DFCIL) for Skeleton Gesture, a new research domain introduced by BOAT-MI[1]. Unlike BOAT-MI which involves data synthesis via model inversion and retraining of the whole model, we introduce a new model-updating paradigm called Synthetic Feature Replay (SFR). SFR confines the synthesis process to the embedding domain by modeling the class feature space as a multivariate normal distribution defined by a prototype (mean and covariance). This approach allows us to directly sample synthetic features to recap old class knowledge and facilitate the learning of new classes.
@@
Based on this observation, our proposed SFR circumvents the need for challenging synthetic data generation, which is typically used in existing DFCIL methods. Our approach is not only faster and simpler, but also significantly outperforms current SOTA methods across all tested datasets.
Additionally, we examine the effect of shot size within the Few Shot Class Incremental Learning (FSCIL) context and found that sampling synthetic features for new classes combined with a carefully designed filtering mechanism can alleviate performance degradation caused by limited data availability.
@@ 2 Related Work
FSCIL [41] is a setup that imposes additional restrictions on the sample size of incremental classes beyond those in DFCIL. One strategy widely used in few-shot class incremental learning is to keep the backbone encoder frozen and update only the classifier to adapt to new classes. These methods focus either on creating better pre-trained base models [55, 40, 56], or designing novel classifiers [52]. Constructing a class prototype [27, 14, 3, 21], originally proposed by ProtoNet [39], primarily used for a few-shot learning, has proved useful in creating the classifier for FSCIL as well [57, 45]. Our approach also adopts such a frozen encoder strategy, but instead of directly using class prototypes to build the classifier, we sample synthetic features from class prototypes and train the classifier along with the sampled features.
@@
Using skeletal data, instead of raw image frames, for gesture recognition shows great value for its reduced data complexity and therefore effective processing and privacy protection. Deep learning approaches such as CNN-LSTM [31], GCNs[49, 7, 23], transformers [6, 36, 54, 44], typically show superior performance for skeleton-based gesture recognition. ST-GCN[49] is a popular GCN-based model that can automatically learn both spatial and temporal patterns. DG-STA [6] is a transformer-based model that applies multi-head attention to spatial and temporal edges, respectively, for hand gesture recognition.
Class incremental learning for skeleton data, especially in the data-free setting, is still a valuable but underexplored domain. To the best of our knowledge, BOAT-MI [1] represents pioneering work in investigating data-free class incremental learning of gestures by developing a model inversion-based approach to generate synthetic data and replay them during training new classes. Therefore, we follow their setup to run our experiments and benchmark against them.
@@ 3 Preliminaries
Instantaneous Forgetting Measure (IFM).
The central dilemma in class incremental learning revolves around balancing the preservation of knowledge from old tasks with the ongoing learning of new tasks. However, many existing studies[26, 52, 56, 8] tend to over-prioritize overall accuracy as the primary evaluation metric, often overlooking the imbalance between the accuracy of the base classes and the incremental classes.
@@
where represents the local accuracy solely over the novel classes introduced during the current step, while signifies the global accuracy over all classes encountered up to that stage.
4 What to Replay? Embeddings or Data
Synthetic replay-based DFCIL approaches[8, 37, 51, 13, 1] typically involve training an additional generator or using model inversion to create fake data. However, generating those fake data poses significant challenges, as pre-training a generator during the previous step can be cumbersome, and the quality of data generated through model inversion may suffer, especially for skeletal gesture movements that entail both spatial and temporal changes.
Thus, is there a way to bypass the challenging data synthesis step? Instead of traditional replay based DFCIL approaches which focus on data synthesis, we proposed a simple but effective paradigm that directly synthesize fake features from the embedding space to replay the old class and facilitate the training of the new class, when the samples of the newly class are limited, as illustrated in Figure 1. The effectiveness of our approach is rooted in a key observation.
@@ 4 What to Replay? Embeddings or Data
However, we found that using a frozen encoder is particularly well suited for skeleton modality, as our experiments unveiled an interesting discovery: models trained on skeletal-based gesture movement data can automatically generalize well to new, unseen classes without additional training. This enables the formation of compact feature spaces that establish distinguishable boundaries from old and new classes. A simple experiment on Shrec-2017[9], as shown in Figure 3(b), illustrates how within-class features cluster together, forming compact class feature spaces with minimal overlap between different classes. This phenomenon appears to be specific to skeletal data (or potentially other structured modalities). Taking a hand gesture as an example, regardless of the specific movement (e.g., pinch or swipe), all data samples are fundamentally hand motions, and the structure should always adhere to the physical appearance of a hand. Thus, treating all data as a single class with various subclasses, explainable by semantic relationships, seems reasonable. Previous studies [30, 10, 22, 12, 32, 29] have shown that both large image and language models demonstrated a strong relationship between feature space and semantic information, especially within a large category. Furthermore, an observation that supports this claim is the significantly higher DFCIL performance on CUB200 (lower SAD and IFM in the last session) compared to two other image datasets, as shown in Figure 2. This can possibly be attributed to the fine-grain nature of the CUB200 which contains 200 subcategories of a single large class - bird. The skeletal encoder, even trained with very limited data, appears to be capable of capturing such a feature relationship. This observation forms the basis for our feature sampling approach.
Synthetic Feature Sampling.
First, the model is trained with initial subset of dataset . Then, the model’s feature encoder is kept frozen for all incremental sessions and only the classifier is continuously updated. Prior to training the current classifier , only the class prototypes for all seen classes before the current incremental step are saved and kept for future use as well.
The class prototypes can be described by the mean and covariance , to provide an accurate approximation of class feature space. Then, we can directly synthesize embeddings from the prototypes by making the assumption that each class feature space follows a multivariate Gaussian distribution (we further investigate the correctness and feasibility of the Gaussian assumption in the appendix). The sampling processing can be formulated as:
@@
where is the synthetic feature space of class sampled from the distribution, is the label space.
@@
Considering that feature representations often reside in high-dimensional spaces, a potential issue arises with the dimensionality-collapse problem. This arises when the computed class covariance fails to be positive semi-definite, contradicting the assumption of a normal distribution. One proposed solution is to employ singular value decomposition (SVD) to reduce the dimensionality of the feature space. Subsequently, sampling can be conducted in this reduced feature space, followed by reverting the sampled features back to their original dimension size, as elaborated in Alg. 1.
After the sampling process is complete, we can obtain the real features of classes newly introduced in the current incremental session by simply feeding the data into the frozen encoder , along with the fake features sampled from saved old class prototypes. These real and fake features are then combined to train the classifier .
@@ 4.1 Effect of Supervised Contrastive Learning
Model trained with constraive learning are proven to generate more robust and generailzable representation, making it effective for multiple downstream tasks, and have a superior transfer ability than its cross-entropy counterpart [18]. Some previous work has shown its effectiveness in continual learning studies [11, 40].
Tables 1 and 2 show the performance when the base model is trained with supervised contrastive loss (Supcon) and cross entropy. For Shrec-2017, which is a small dataset, the model trained with Supcon indicates a great performance for both overall accuracy and more significantly for the IFM score, which is half compared to model trained with cross entropy. However, for the much larger Ego-Gesture3D, the significance of contrastive learning is not significant.
@@ 4.3 SOTA Comparison
Efficiency and Privacy.
In addition to its superior performance, feature synthesis offers the significant advantage of low computational cost. This approach only requires simple sampling from the embedding space, which typically takes just a few seconds or less. In contrast, other replay-based methods often involve training generators and creating pseudo-data, both of which are computationally expensive. For instance, BOAT-MI relies on the model inversion process, which can take hours just to generate fake samples. Furthermore, synthetic data might still reveal certain characteristics of the real training samples, which may conflict with the goal of data-free methods aimed at enhancing privacy. SFR also safeguards privacy by confining the synthesis process within the embedding domains.
Method Task 0 Task 1 Task 2 Task 3 Task 4 Task 5 Task 6 Mean (Task G IFM G IFM G IFM G IFM G IFM G IFM G IFM BOAT-MI (Full-Shot) 78.1 75.3 4.6 70.9 7.3 64.0 15.1 58.6 16.3 52.3 26.8 45.8 32.2 61.2 17.1 TEEN (5-Shot) 88.7 83.4 30.5 76.1 40.6 71.3 40.8 72.4 23.4 68.0 25.5 63.8 27.9 74.8 27.0 SFR (Full-Shot) 89.4 88.4 3.3 87.3 2.1 86.0 0.5 85.3 0.2 83.0 0.2 82.3 1.9 85.9 1.2 SFR (10-Shot) 89.4 88.0 6.2 86.3 5.8 81.8 12.1 85.4 1.2 78.8 6.9 78.9 6.2 84.1 5.5 SFR (5-Shot) 89.4 85.9 20.7 82.8 15.8 79.9 13.2 78.5 9.8 75.9 11.1 74.4 10.1 81.0 11.5 SFR-A (5-Shot) 89.4 87.2 10.0 84.2 10.5 79.2 9.7 79.8 5.9 75.0 10.2 74.4 5.9 81.3 7.5
Method Task 0 Task 1 Task 2 Task 3 Task 4 Task 5 Task 6 Mean (Task G IFM G IFM G IFM G IFM G IFM G IFM G IFM BOAT-MI (Full-Shot) 78.1 75.3 4.6 70.9 7.3 64.0 15.1 58.6 16.3 52.3 26.8 45.8 32.2 61.2 17.1 TEEN (5-Shot) 77.1 74.6 28.9 72.5 26.9 70.1 27.0 69.0 22.7 66.5 24.9 64.6 24.7 70.6 22.2 SFR (Full-Shot) 75.2 75.6 2.5 73.1 3.0 72.7 1.7 72.7 2.4 70.2 2.2 69.5 3.2 72.7 2.1 SFR (10-Shot) 75.2 73.6 14.5 71.8 9.0 68.9 11.5 66.5 8.1 66.3 6.5 63.0 13.6 69.4 9.0 SFR (5-Shot) 75.2 72.1 15.8 68.4 25.2 65.1 18.6 64.4 15.2 60.8 19.3 58.9 20.2 66.4 16.3 SFR-A (5-Shot) 75.2 72.3 10.8 69.1 14.2 66.6 15.7 66.4 12.5 62.6 11.9 58.3 18.0 67.2 11.9 *SFR-A denotes feature augmentation is enabled while SFR is disabled.